# Viscosity of polymer melts using non-affine theory based on vibrational modes

### v1

@@ I Introduction
The structure and dynamics of polymers are crucial for the development of different chemical and biological applications [1]. Over the past decades, extensive studies on both naturally occurring and synthetic polymers, such as proteins and chemical compounds, have advanced our understanding in these areas. Polymer dynamics depends on the structural arrangement, length scales, and topological constraints imposed by the surrounding chains. Early work, such as single-chain studies by de Gennes [2], laid the foundation to the understanding of these dynamics. Over the years, various groups have investigated the dynamical and rheological properties of polymer melts using theoretical, simulation, and experimental approaches, revealing phenomena such as contour length fluctuations [3, 4], constraint release [5, 6], collapse kinetics [7, 8, 9], confinement effects [10], and phase behavior [11, 12]. Recently, significant attention has been directed to the study of the rheological properties of entangled and ring polymer melts [13, 14, 15, 16, 17, 18]. Polymer melts, with their amorphous properties, serve as effective models for supercooled liquids [19, 20]. Although tube theory and reptation concepts have proven invaluable in obtaining a deeper molecular level understanding of polymer dynamics [1], for many applications and materials design it is becoming increasingly crucial to be able to quantitatively predict rheological properties from atomic-scale structure and interactions [21, 22], something that mean-field methods cannot achieve.
@@
Viscosity is a fundamental property for understanding the transport behavior of liquids and supercooled liquids [23, 24, 25, 26]. In supercooled liquids, viscosity increases rapidly upon cooling, making it difficult to measure it near the glass transition temperature [27, 28]. Various methods are available to calculate the viscosity in liquids and amorphous systems, including equilibrium and non-equilibrium approaches [29, 30, 31, 32]. In equilibrium methods, viscosity is derived from pressure or momentum fluctuations, while non-equilibrium methods rely on measuring the mechanical response to external deformation. The Green-Kubo (GK) method [33, 34] is the most widely used equilibrium approach, whereby viscosity is calculated by integrating the stress autocorrelation function. These correlations decay slowly, and it is difficult to measure smooth long-time tails of such correlations. In general, non-equilibrium methods require very small shear rates to measure Newtonian viscosity, and also extensive averaging due to the highly fluctuating nature of the stress-strain behavior. Consequently, these methods are suitable within a limited temperature range for supercooled liquids. Furthermore, the theoretical approach to derive a microscopic expression for viscosity involves linking particle probability distributions and interatomic potentials using the Born-Green approximation [35]. By employing a hole model of the liquid state, an analytical relationship between viscosity and the thermodynamic properties of dense liquids has been established through the determination of an equation of state [36, 37]. Other approaches, such as activation rate theory [38, 39, 40], relate viscosity or relaxation times to the hopping rates of particles as they move across the cage formed by their nearest neighbors.
@@
There have been attempts to relate the frequency-dependent shear modulus of viscoelastic liquids to viscosity. One such phenomenological framework is based on Maxwell’s model [41], and several proposals have explored the connection between the viscoelastic modulus and transport properties, such as viscosity, in supercooled liquids [29, 30, 31, 42, 43, 44]. However, a significant theoretical advancement in this direction is the development of non-affine lattice dynamics (NALD), where the equation of motion for the non-affine displacement of a tagged particle in a disordered medium is explicitly solved, providing deeper insight into the microscopic mechanisms governing viscosity. The NALD approach relies on specific microscopic inputs, primarily the nature of the potential energy surface and the Hessian matrix of the system, to predict the frequency-dependent mechanical response of the system [26].
@@
At its core, the NALD framework is based on the fundamental concept of non-affine displacements [45, 46]. When a material sample undergoes deformation, each atom tends to follow the applied strain, and the affine displacement is defined as the atomic displacement from its original position in the undeformed sample to the position prescribed by the macroscopic strain. In this (affine) position, due to the lack of centrosymmetry, the atom is not at mechanical equilibrium and must undergo an additional displacement, which is the non-affine displacement [47, 21]. The force acting on the atom in the affine position, which triggers the non-affine displacement, is called the affine force or affine force field [26]. The solution, expressed in terms of the complex viscoelastic modulus as a function of external frequency, is then obtained after performing projection of the atomic displacements onto the eigenvectors of the Hessian matrix, and upon Fourier transformation. Finally, the viscosity is computed using the viscoelastic modulus within the framework of non-affine response theory [48].
@@
In this paper, we investigate the microscopic origin of viscosity in supercooled liquids from first principles using non-affine response theory based on the NALD framework. We study a coarse-grained polymer melt in the supercooled regime by employing molecular dynamics (MD) simulations to generate the configurations [49, 50]. These configurations are then used to calculate the vibrational density of states (vDOS) and affine force field correlators, which are subsequently used to compute the complex shear modulus. The viscosity of the system in the zero-frequency limit is directly obtained from the loss modulus following the theory of Ref. [48]. The viscosity in the supercooled regime and its temperature dependence is accurately captured by this NALD approach. We also calculate the viscosity using the GK method and non-equilibrium shear MD simulations over the accessible temperature range for supercooled liquids. By comparing the viscosity results obtained from different methods in both the liquid and supercooled regimes, we show that the NALD approach reliably predicts the viscosity, particularly in the supercooled regime where experimental measurements and simulations based on the conventional methods are challenging.
@@ II Model and methods
We have studied a coarse-grained polymer, popularly known as Kremer-Grest model [51], where the polymer chain has beads (monomers) connected linearly via non-breakable covalent bonds, represented by finite extensible nonlinear elastic (FENE) potential [49]. Such a potential has the form $U_{\rm FENE}=-0.5KR_{0}^{2}\ln[1-(r/R_{0})^{2}]$. In this model, the monomers interact pairwise with all other monomers via a Lennard-Jones (LJ) interaction $U_{\rm LJ}=V(r)-V(r_{c})$, where $V(r)=4\epsilon[(\sigma/r)^{12}-(\sigma/r)^{6}]$, for $r<\sigma$, and $r$ is the inter-monomer separation. All the measurements are performed in standard reduced LJ units. In this study, we set the parameters as follows: $K=30,R_{0}=1.5,\epsilon=1,\sigma=1,r_{c}=2.5$. The mass of each monomer is set to unity. We have only these two types of interaction present in the system. Since there are no bending potentials in the form of three monomer angle, four monomer dihedral, etc., the polymer chain is fully flexible. We have simulated $M=100$ polymer chains, each consisting of $50$ identical beads. This makes a system of $N=5000$ particles.
@@
Large-scale molecular dynamics simulations are performed using LAMMPS [52], where equations of motion are integrated via velocity-Verlet algorithm with timestep $\Delta t=0.005$ under periodic boundary conditions. To prepare an equilibrated configuration in the supercooled state, we start with a random configuration of a polymer chain, prepared via self-avoiding random walk. These random configurations are first heated at a sufficiently high temperature and zero pressure, and then quenched to the target temperature and equilibrated for a sufficiently long time, maintaining the zero pressure. Langevin thermostat and Nosè-Hoover barostat, both having unit dissipation timescale, are employed to maintain the temperature and pressure of the system. For the system we have used in our study, the glass transition temperature is[50, 49] $T_{g}\simeq 0.4$. We have generated at least $30$ independent configurations at each temperature in the range $\left[0.6,1.8\right]$.
@@
In this work, we have also studied the response of the system to shear deformation. Shear is imposed by deforming the $xy$-plane in the $x$-direction [53, 54], at different shear rates $\dot{\gamma}$ ranging between $10^{-4}$ and $5\times 10^{-3}$. In the presence of shear, Lees-Edwards boundary conditions are utilized to take into account the deformed boundary [55]. Furthermore, during deformation, the temperature of the system is maintained using dissipative particle dynamics (DPD) thermostat [56, 57] without any external pressure control. Shear stress $\sigma_{xy}$ is measured using the Irving-Kirkwood definition: $\sigma_{xy}=\langle\frac{1}{V}\sum_{i,j}f^{x}_{i,j}r_{i,j}^{y}\rangle$, where $f^{x}_{i,j}$ is the $x$-component of the force acting on particle $i$ due to $j$, $r_{i,j}^{y}$ is the $y$-component of the inter-particle separation between $i$ and $j$, and $V$ is the volume of the system. We study the evolution of shear-stress $\sigma_{xy}$ as a function of deformation in the form of shear-strain $\gamma$ at different $\dot{\gamma}$.
@@
For the normal mode analysis, we compute the dynamical matrix or Hessian ${\bf H}$, using the method of finite-differences. Instantaneous configurations at different temperatures are utilized for these calculations. For our three-dimensional system of $N$ particles, the Hessian is a $3N\times 3N$ matrix. The elements of the matrix are defined as ${\bf H}_{ij}^{\alpha\beta}=\frac{1}{\sqrt{m_{i}m_{j}}}\frac{\delta^{2}U}{\delta r_{i,\alpha}\delta r_{j,\beta}}$, where $i,j$ are particle index, $\alpha,\beta$ are Cartesian directions, and $U$ is the system potential energy. We diagonalize the matrix using the LAPACK package implemented in Intel MKL library [58]. This results in $3N$ eigenvalues $\lambda_{l}$ and the same number of eigenvectors ${\bf e}_{l}$, $l={1,2,3,...,3N}$. An eigenfrequency $\omega$ is defined corresponding to each eigenvalue, $\omega_{l}=\sqrt{\lambda_{l}}$. Since we study finite temperature configurations, we get a lot of negative eigenvalues after diagonalization, which result in imaginary eigenfrequencies [59, 60], also known as instantaneous normal modes (INMs) [61, 62, 63, 64].
@@ II.1 Non-affine lattice dynamics approach to viscosity
In an ideal elastic solid, when deformation is applied, particles move toward their affine positions and the local inversion symmetry ensures that forces from neighboring particles cancel out, leaving the particle in mechanical equilibrium at the affine position. However, the basis of non-affine lattice dynamics (NALD) lies in recognizing that, in disordered (amorphous) solids, the standard affine approximation of lattice dynamics breaks down. In such systems due to the absence of centrosymmetry, the applied deformation generates non-zero net forces on atoms, leading to additional non-affine displacements. This process requires internal work to displace particles from their affine to non-affine positions, which contributes negatively to the free energy of deformation. The equation of motion for a tagged particle $i$ of mass $m$ within the framework of the generalized Langevin equation [65, 59, 66] reads as
@@
Here, $\textbf{r}_{i}$ represents the position of the $i$th particle, $\nu$ is the friction coefficient that characterizes energy dissipation due to viscous damping at the atomic scale[66], the third term (with ${\bf H}_{ij}$ the Hessian matrix) represents the sum of the restoring forces arising from interactions with neighboring particles, and the term containing $\Xi$ represents the affine force induced by the applied shear strain $\gamma$.
@@
where, $G_{A}$ is the modulus due to affine part, $V$ is volume of system, $\omega_{k}$ are the normal modes of system and $\Gamma(\omega_{k})=\tilde{\Xi}^{2}_{k}$ is the affine force field correlator [45, 67, 59]. Equation 2 can be expressed in a continuous form by using the normalized vibrational density of states $g(\omega)$
@@
By separating the real and imaginary parts of Eq. (3), we obtain the storage modulus ($G^{\prime}$) and the loss modulus ($G^{\prime\prime}$) [48, 59],
@@
The vibrational density of states (vDOS), $g(\omega)$, is defined using the normal modes obtained after the diagonalization of the Hessian matrix constructed with the instantaneous configuration of the system [26, 68]: $g(\omega)=\frac{1}{3N-3}\sum_{i}\delta(\omega-\omega_{i})$, where we discard three Goldstone modes appearing due to rigid-body translations. Such vDOS can be derived analytically [69] or computed numerically [64, 50], and, also, can be extracted from experimental data [70, 71]. The frequency-dependent affine force field correlator, $\Gamma(\omega)$, is calculated as the square of the projection of the force field $\Xi$ onto the eigenvectors corresponding to the normal modes. The affine force field, which drives non-affine motion in the sheared system, is calculated as the net force acting on each atom after the system is affinely deformed by a very small shear strain $\eta=10^{-6}$. Specifically, $\Xi_{i}=\Delta f_{i}/\eta$, where $\Delta f_{i}$ represents the net force on atom $i$ relative to its position in the undeformed state. The analytical estimate of $\Gamma(\omega)$ at low eigenfrequencies, derived in Ref. [46], is $\Gamma(\omega)\sim\omega^{2}$, and has been proved to be a reasonable approximation for amorphous solids [72, 73]. To calculate the affine modulus $G_{A}$, the system is deformed affinely with a small shear strain $\eta=10^{-6}$ and the corresponding shear stress $\sigma$ is measured. Then, the affine modulus is calculated as $G_{A}=\sigma/\eta$.
The non-affine response theory, developed from first principles, establishes a relationship between the shear viscosity and the loss modulus $G^{\prime\prime}$, given by [48]
@@
Throughout this study, we use a fixed value of ${\nu}=1$ in the above equations, which matches exactly the value implemented in the Langevin thermostat of the MD simulations [50]. Hence, there are no adjustable parameters in our calculations. One could consider varying the memory friction parameter $\nu$ and observing its effect on the viscoelastic modulus [59] and thermodynamic properties. In a Markovian process, friction is treated as a constant, as implemented in molecular simulations using the Langevin thermostat. However, in real materials such as metallic glasses or experimental systems, friction is better described within a non-Markovian framework, where the friction kernel is time-dependent or history-dependent [74, 75]. The friction kernel can be computed using the fluctuation-dissipation theorem (FDT), allowing us to determine the memory kernel in the frequency domain. This frequency-dependent memory kernel can then be incorporated into the viscoelastic modulus, providing a more accurate description of real materials and experimental conditions, which is a whole task in its own right that we leave for future studies.

### v3

@@ I Introduction
The structure and dynamics of polymers are crucial for the development of different chemical and biological applications Doi et al. (1988). Over the past decades, extensive studies on both naturally occurring and synthetic polymers, such as proteins and chemical compounds, have advanced our understanding in these areas. Polymer dynamics depends on the structural arrangement, length scales, and topological constraints imposed by the surrounding chains. Early work, such as single-chain studies by de Gennes De Gennes (1979), laid the foundation to the understanding of these dynamics. Over the years, various groups have investigated the dynamical and rheological properties of polymer melts using theoretical, simulation, and experimental approaches, revealing phenomena such as contour length fluctuations Doi (1983); Milner and McLeish (1998), constraint release Rubinstein and Colby (1988); Shivokhin et al. (2014), collapse kinetics Byrne et al. (1995); Kikuchi et al. (2005); Chauhan and Singh (2022), confinement effects McLeish (2002), and phase behavior Poier et al. (2015); Nikoubashman (2021). Recently, significant attention has been directed to the study of the rheological properties of entangled and ring polymer melts Pasquino et al. (0 15); Liebetreu and Likos (2020); Parisi et al. (2021); Datta et al. (2023); Roy et al. (2022); Peponaki et al. (2024). Polymer melts, with their amorphous properties, serve as effective models for supercooled liquids Dudowicz et al. (2005); Vaibhav and Dutta (2024). Although tube theory and reptation concepts have proven invaluable in obtaining a deeper molecular level understanding of polymer dynamics Doi et al. (1988), for many applications and materials design it is becoming increasingly crucial to be able to quantitatively predict rheological properties from atomic-scale structure and interactions Theodorou and Suter (1986); Vogiatzis et al. (2020), something that mean-field methods cannot achieve.
@@
Viscosity is a fundamental property for understanding the transport behavior of liquids and supercooled liquids Viswanath et al. (2007); Huang et al. (2022, 2025a); Trachenko (2023); Zaccone (2023a); Hopper et al. (2024). In supercooled liquids, viscosity increases rapidly upon cooling, making it difficult to measure it near the glass transition temperature Ediger et al. (1996); Debenedetti and Stillinger (2001). Various methods are available to calculate the viscosity in liquids and amorphous systems, including equilibrium and non-equilibrium approaches Sen et al. (2005); Kröger et al. (1993); Hess (2002); Evans and Morriss (2008). In equilibrium methods, viscosity is derived from pressure or momentum fluctuations, while non-equilibrium methods rely on measuring the mechanical response to external deformation. The Green-Kubo (GK) method Green (1954); Kubo (1957) is the most widely used equilibrium approach, whereby viscosity is calculated by integrating the stress autocorrelation function. These correlations decay slowly, and it is difficult to measure smooth long-time tails of such correlations. In general, non-equilibrium methods require very small shear rates to measure Newtonian viscosity, and also extensive averaging due to the highly fluctuating nature of the stress-strain behavior. Consequently, these methods are suitable within a limited temperature range for supercooled liquids. Furthermore, the theoretical approach to derive a microscopic expression for viscosity involves linking particle probability distributions and interatomic potentials using the Born-Green approximation Born and Green (1947). By employing a hole model of the liquid state, an analytical relationship between viscosity and the thermodynamic properties of dense liquids has been established through the determination of an equation of state Sanchez (1974); Locati et al. (1999). Other approaches, such as activation rate theory Glasstone et al. (1941); Mauro et al. (2009); Dyre (2006); Schweizer (2005); Singh et al. (2021); Singh and Singh (2023), relate viscosity or relaxation times to the hopping rates of particles as they move across the cage formed by their nearest neighbors.
@@
There have been attempts to relate the frequency-dependent shear modulus of viscoelastic liquids to viscosity. One such phenomenological framework is based on Maxwell’s model Frenkl (1955), and several proposals have explored the connection between the viscoelastic modulus and transport properties, such as viscosity, in supercooled liquids Sen et al. (2005); Kröger et al. (1993); Hess (2002); Marrucci and Ianniruberto (2000); Peluso (2024); Huang et al. (2025b). However, a significant theoretical advancement in this direction is the development of non-affine lattice dynamics (NALD), where the equation of motion for the non-affine displacement of a tagged particle in a disordered medium is explicitly solved, providing deeper insight into the microscopic mechanisms governing viscosity. The NALD approach relies on specific microscopic inputs, primarily the nature of the potential energy surface and the Hessian matrix of the system, to predict the frequency-dependent mechanical response of the system Zaccone (2023a).
@@
At its core, the NALD framework is based on the fundamental concept of non-affine displacements Lemaître and Maloney (2006); Zaccone and Scossa-Romano (2011). When a material sample undergoes deformation, each atom tends to follow the applied strain, and the affine displacement is defined as the atomic displacement from its original position in the undeformed sample to the position prescribed by the macroscopic strain. In this (affine) position, due to the lack of centrosymmetry, the atom is not at mechanical equilibrium and must undergo an additional displacement, which is the non-affine displacement Slonczewski and Thomas (1970); Theodorou and Suter (1986). The force acting on the atom in the affine position, which triggers the non-affine displacement, is called the affine force or affine force field Zaccone (2023a). The solution, expressed in terms of the complex viscoelastic modulus as a function of external frequency, is then obtained after performing projection of the atomic displacements onto the eigenvectors of the Hessian matrix, and upon Fourier transformation. Finally, the viscosity is computed using the viscoelastic modulus within the framework of non-affine response theory Zaccone (2023b).
@@
In this paper, we investigate the microscopic origin of viscosity in supercooled liquids from first principles using non-affine response theory based on the NALD framework. We study a coarse-grained polymer melt in the supercooled regime by employing molecular dynamics (MD) simulations to generate the configurations Kremer and Grest (1990); Palyulin et al. (2018). These configurations are then used to calculate the vibrational density of states (vDOS) and affine force field correlators, which are subsequently used to compute the complex shear modulus. The viscosity of the system in the zero-frequency limit is directly obtained from the loss modulus following the theory of Ref. Zaccone (2023b). The viscosity in the supercooled regime and its temperature dependence is accurately captured by this NALD approach. We also calculate the viscosity using the GK method and non-equilibrium shear MD simulations over the accessible temperature range for supercooled liquids. By comparing the viscosity results obtained from different methods in both the liquid and supercooled regimes, we show that the NALD approach reliably predicts the viscosity, particularly in the supercooled regime where experimental measurements and simulations based on the conventional methods are challenging.
@@ II Model and methods
We have studied a coarse-grained polymer, popularly known as Kremer-Grest model Grest and Kremer (1986), where the polymer chain has beads (monomers) connected linearly via non-breakable covalent bonds, represented by finite extensible nonlinear elastic (FENE) potential Kremer and Grest (1990). Such a potential has the form $U_{\rm FENE}=-0.5KR_{0}^{2}\ln[1-(r/R_{0})^{2}]$. In this model, the monomers interact pairwise with all other monomers via a Lennard-Jones (LJ) interaction $U_{\rm LJ}=V(r)-V(r_{c})$, where $V(r)=4\epsilon[(\sigma/r)^{12}-(\sigma/r)^{6}]$, for $r<\sigma$, and $r$ is the inter-monomer separation. All the measurements are performed in standard reduced LJ units. In this study, we set the parameters as follows: $K=30,R_{0}=1.5,\epsilon=1,\sigma=1,r_{c}=2.5$. The mass of each monomer is set to unity. We have only these two types of interaction present in the system. Since there are no bending potentials in the form of three monomer angle, four monomer dihedral, etc., the polymer chain is fully flexible. We have simulated $M=100$ polymer chains, each consisting of $50$ identical beads. This makes a system of $N=5000$ particles.
@@
Large-scale molecular dynamics simulations are performed using LAMMPS Plimpton (1995), where equations of motion are integrated via velocity-Verlet algorithm with timestep $\Delta t=0.005$ under periodic boundary conditions. To prepare an equilibrated configuration in the supercooled state, we start with a random configuration of a polymer chain, prepared via self-avoiding random walk. These random configurations are first heated at a sufficiently high temperature and zero pressure, and then quenched to the target temperature and equilibrated, maintaining the zero pressure. Langevin thermostat and Nosè-Hoover barostat, both having unit dissipation timescale, are employed to maintain the temperature and pressure of the system. For the system we have used in our study, the glass transition temperature isPalyulin et al. (2018); Kremer and Grest (1990) $T_{g}\simeq 0.4$. We have generated at least $30$ independent configurations at each temperature in the range $\left[0.6,1.8\right]$.
@@
In this work, we have also studied the response of the system to shear deformation. Shear is imposed by deforming the $xy$-plane in the $x$-direction Vaibhav et al. (2022, 2023), at different shear rates $\dot{\gamma}$ ranging between $10^{-4}$ and $5\times 10^{-3}$. In the presence of shear, Lees-Edwards boundary conditions are utilized to take into account the deformed boundary Allen and Tildesley (2017). Furthermore, during deformation, the temperature of the system is maintained using dissipative particle dynamics thermostat Soddemann et al. (2003); Vaibhav and Chaudhuri (2021) without any external pressure control. Shear stress $\sigma_{xy}$ is measured using the virial definition: $\sigma_{xy}=\langle\frac{1}{V}\sum_{i,j}f^{x}_{i,j}r_{i,j}^{y}\rangle$, where $f^{x}_{i,j}$ is the $x$-component of the force acting on particle $i$ due to $j$, $r_{i,j}^{y}$ is the $y$-component of the inter-particle separation between $i$ and $j$, and $V$ is the volume of the system.Thompson et al. (2009) We study the evolution of shear-stress $\sigma_{xy}$ as a function of deformation in the form of shear-strain $\gamma$ at different $\dot{\gamma}$.
@@
For the normal mode analysis, we compute the dynamical matrix or Hessian ${\bf H}$, using the method of finite-differences. Instantaneous configurations at different temperatures are utilized for these calculations. For our three-dimensional system of $N$ particles, the Hessian is a $3N\times 3N$ matrix. The elements of the matrix are defined as ${\bf H}_{ij}^{\alpha\beta}=\frac{1}{\sqrt{m_{i}m_{j}}}\frac{\delta^{2}U}{\delta r_{i,\alpha}\delta r_{j,\beta}}$, where $i,j$ are particle index, $\alpha,\beta$ are Cartesian directions, and $U$ is the system potential energy. We diagonalize the matrix using the LAPACK package implemented in Intel MKL library Anderson et al. (1999). This results in $3N$ eigenvalues $\lambda_{l}$ and the same number of eigenvectors ${\bf e}_{l}$, $l={1,2,3,...,3N}$. An eigenfrequency $\omega$ is defined corresponding to each eigenvalue, $\omega_{l}=\sqrt{\lambda_{l}}$. Since we study finite temperature configurations, negative eigenvalues are present after diagonalization, which result in imaginary eigenfrequencies Palyulin et al. (2018); Vaibhav et al. (2024), also known as instantaneous normal modes (INMs) Keyes (1994); Stratt (1995); Bembenek and Laird (1995); Zhang et al. (2019).
@@ II.1 Non-affine lattice dynamics approach to viscosity
In an ideal elastic solid, when deformation is applied, particles move toward their affine positions and the local inversion symmetry ensures that forces from neighboring particles cancel out, leaving the particle in mechanical equilibrium at the affine position. However, the basis of non-affine lattice dynamics (NALD) lies in recognizing that, in disordered (amorphous) solids, the standard affine approximation of lattice dynamics breaks down. In such systems due to the absence of centrosymmetry, the applied deformation generates non-zero net forces on atoms, leading to additional non-affine displacements. This process requires internal work to displace particles from their affine to non-affine positions, which contributes negatively to the free energy of deformation. The equation of motion for a tagged particle $i$ of mass $m$ within the framework of the generalized Langevin equation Cui et al. (2017); Palyulin et al. (2018); Zwanzig (2001) reads as
@@
Here, $\textbf{r}_{i}$ represents the position of the $i$th particle, $\nu$ is the friction coefficient that characterizes energy dissipation due to viscous damping at the atomic scaleZwanzig (2001), the third term (with ${\bf H}_{ij}$ the Hessian matrix) represents the sum of the restoring forces arising from interactions with neighboring particles, and the term containing $\Xi$ represents the affine force induced by the applied shear strain $\gamma$.
@@
where, $G_{A}$ is the modulus due to affine part, $V$ is volume of system, $\omega_{k}$ are the normal modes of system and $\Gamma(\omega_{k})=\tilde{\Xi}^{2}_{k}$ is the affine force field correlator Lemaître and Maloney (2006); Milkus et al. (2018); Palyulin et al. (2018). Equation 2 can be expressed in a continuous form by using the normalized vibrational density of states $g(\omega)$
@@
By separating the real and imaginary parts of Eq. (3), we obtain the storage modulus ($G^{\prime}$) and the loss modulus ($G^{\prime\prime}$) Zaccone (2023b); Palyulin et al. (2018),
@@
The vibrational density of states (vDOS), $g(\omega)$, is defined using the normal modes obtained after the diagonalization of the Hessian matrix constructed with the instantaneous configuration of the system Zaccone (2023a); Hara et al. (2025): $g(\omega)=\frac{1}{3N-3}\sum_{i}\delta(\omega-\omega_{i})$, where we discard three Goldstone modes appearing due to rigid-body translations. Such vDOS can be derived analytically Zaccone and Baggioli (2021) or computed numerically Zhang et al. (2019); Palyulin et al. (2018), and can also be extracted from experimental data Stamper et al. (2022); Vaibhav et al. (2025). The frequency-dependent affine force field correlator, $\Gamma(\omega)$, is calculated as the square of the projection of the force field $\Xi$ onto the eigenvectors corresponding to the normal modes. The affine force field, which drives non-affine motion in the sheared system, is calculated as the net force acting on each atom after the system is affinely deformed by a very small shear strain $\eta=10^{-6}$. Specifically, $\Xi_{i}=\Delta f_{i}/\eta$, where $\Delta f_{i}$ represents the net force on atom $i$ relative to its position in the undeformed state. The analytical estimate of $\Gamma(\omega)$ at low eigenfrequencies, derived in Ref. Zaccone and Scossa-Romano (2011), is $\Gamma(\omega)\sim\omega^{2}$, and has been proved to be a reasonable approximation for amorphous solids Grießer and Pastewka (2024); Flenner and Szamel (2024). To calculate the affine modulus $G_{A}$, the system is deformed affinely with a small shear strain $\eta=10^{-6}$ and the corresponding shear stress $\sigma$ is measured. Then, the affine modulus is calculated as $G_{A}=\sigma/\eta$.
The non-affine response theory, developed from first principles, establishes a relationship between the shear viscosity and the loss modulus $G^{\prime\prime}$, given by Zaccone (2023b)
@@
Throughout this study, we use a fixed value of ${\nu}=1$ in the above equations, which matches exactly the value implemented in the Langevin thermostat of the MD simulations Palyulin et al. (2018). Hence, there are no adjustable parameters in our calculations. One could consider varying the memory friction parameter $\nu$ and observing its effect on the viscoelastic modulus Palyulin et al. (2018) and thermodynamic properties. In a Markovian process, friction is treated as a constant, as implemented in molecular simulations using the Langevin thermostat. However, in real materials such as metallic glasses or experimental systems, friction is better described within a non-Markovian framework, where the friction kernel is time-dependent or history-dependent Klippenstein et al. (2021); Milster et al. (2024). The friction kernel can be computed using the fluctuation-dissipation theorem (FDT), allowing us to determine the memory kernel in the frequency domain. This frequency-dependent memory kernel can then be incorporated into the viscoelastic modulus, providing a more accurate description of real materials and experimental conditions, which is a whole task in its own right that we leave for future studies.